# Yuehao Bai

### v2

@@ Abstract
This paper considers the problem of design-based inference for the average treatment effect in finely stratified experiments. Here, by “design-based” we mean that the only source of uncertainty stems from the randomness in treatment assignment itself; by “finely stratified” we mean units are first stratified into groups of a fixed size $k$ according to baseline covariates and then, within each group, a fixed number $\ell<k$ are assigned uniformly at random to treatment and the remainder to control. In this setting, we first show under mild conditions that inference using the usual difference-in-means estimator requires an estimator of its variance that is at least asymptotically upward-biased. We then present a novel estimator of the variance and show that it is upward-biased; furthermore, the magnitude of the bias depends in a natural way on the quality of the stratification. Importantly, this estimator remains well-defined even in the challenging setting in which $\ell=1$ or $k-\ell=1$. We then compare our estimator with some well-known estimators that have been proposed previously for this case. We first show that, while these estimators are also upward-biased, the magnitude of their bias does not change in the natural way with the quality of stratification. To further discriminate among these different estimators, we then introduce a framework motivated by a thought experiment in which the finite population can be modeled as having been drawn once in an i.i.d. fashion from a well-behaved probability distribution. In this framework, we argue that our estimator dominates the others in terms of limiting bias, which leads to more precise inferences, and that these improvements are strict except under exceptionally strong restrictions on the treatment effects. Finally, we illustrate the practical relevance of our theoretical results through a simulation study, which reveals that our estimator can in fact lead to substantially more precise inferences, especially when the quality of stratification is high.
@@ 1 Introduction
This paper considers the problem of design-based inference on the average treatment effect in finely stratified experiments. Here, by “design-based” we mean that there is no sampling uncertainty and the only randomness stems from the variation in treatment assignment itself; by “finely stratified” we mean that treatment is assigned in the following fashion: units are first stratified into groups of a fixed size $k$ according to baseline covariates and then, within each group, a fixed number $\ell<k$ are assigned uniformly at random to treatment and the remainder to control. A prominent special case of this framework is a matched pairs design, in which $k=2$ and $\ell=1$.
@@
We then compare our variance estimator with two well-known variance estimators that have been proposed previously for the case in which $\ell=1$ or $k-\ell=1$. The first estimator, proposed by Imai (2008), is based on the sample variance obtained by viewing the difference-in-means estimates for each stratum as independent observations. The second estimator, proposed by Fogarty (2018a), modifies the previous estimator by partialling out the average covariate values in each stratum. Existing results from these papers show that both estimators are also upward-biased for the variance of the difference-in-means estimator but, in contrast to our estimator, their biases do not depend directly on the quality of the stratification. To further discriminate among these estimators and ours, we introduce a framework motivated by a thought experiment in which the finite population can be modeled as being drawn once in an i.i.d. fashion from a well-behaved probability distribution, and the strata are formed in such a way that the covariate values of the units in each stratum, as well as the paired strata, get close in the limit. In this framework, we show that the limiting bias of our estimator is strictly smaller than that of Imai (2008) unless the treatment effects are homogeneous, and strictly smaller than that of Fogarty (2018a) unless the conditional average treatment effect is linear in the covariates. As a result, confidence intervals based on our variance estimator will be strictly shorter than those based on the other variance estimators unless these exceptional restrictions hold.
@@ 2 Setup and Notation
$\hat{\Delta}_{n}^{\dagger}=\frac{1}{n(1)}\sum_{1\leq i\leq n}Y_{i}^{\dagger}D_{i}-\frac{1}{n(0)}Y_{i}^{\dagger}(1-D_{i})~{},$
@@ 3.1 Motivating Result
When (7) holds, we say that $\tilde{V}_{n}$ is upward-biased for $\operatorname{Var}[\hat{\Delta}_{n}]$. Theorem 3.1 thus demonstrates that, under mild regularity conditions, confidence intervals for $\Delta_{n}$ of the form $[\hat{\Delta}_{n}\pm\sqrt{\tilde{V}_{n}}\cdot z_{1-\alpha/2}]$ are valid in the sense of having the right limiting coverage probability, if and only if $\tilde{V}_{n}$ is asymptotically upward-biased for $\operatorname{Var}[\hat{\Delta}_{n}]$. These assumptions include weak restrictions on the finite population (i.e., (2)), a non-degeneracy condition (i.e., (3)), and a requirement that $n\cdot\tilde{V}_{n}$ is consistent for its expectation (i.e., (4)). This last condition will typically hold under some further restrictions on finite population moments; see, e.g., Theorem 4.2 below.
@@ 3.2 Review of Some Upwards Biased Estimators of $\operatorname{Var}[\hat{\Delta}_{n}]$
is the difference in means in the $j$th stratum. This variance estimator serves as the basic scaffolding for many of the recent estimators proposed in the literature on inference in stratified randomized experiments: it is a special case of the small-block variance estimator proposed in Pashley and Miratrix (2021) when all strata are the same size (see also Zhu et al. (2024)), a special case of the pair-cluster variance estimator of De Chaisemartin and Ramirez-Cuellar (2024) in the setting of an individual-level randomized experiment, and a special case of the regression-based estimator due to Fogarty (2018a), which we now describe. Let $R$ be an $m\times L$ matrix for $L<m$ with rank $L$ and $H_{R}=R(R^{\prime}R)^{-1}R^{\prime}$ denote the projection matrix onto its column space. For each $R$, consider the estimator given by
for $\bar{X}_{j,n}=\frac{1}{k}\sum_{i\in\lambda_{j}}X_{i}$ and $\mu_{X,n}=\frac{1}{n}\sum_{1\leq i\leq n}X_{i}$. In particular, Fogarty (2018a) argues that $\hat{V}_{n}^{\rm F}$ can be less conservative than $\hat{V}_{n}^{\rm IM}$ whenever the covariates are predictive of the treatment effects in an appropriate sense.
Note that the bias of $\hat{V}^{\rm IM}_{n}$ is given by (see Imbens and Rubin, 2015; Fogarty, 2018a)
@@
so that $\hat{V}^{\rm IM}_{n}$ is an upward-biased estimator of $\operatorname{Var}[\hat{\Delta}_{n}]$. The bias for $\hat{V}_{n}^{\rm F}$ is (see Fogarty, 2018a)
where $\delta_{n}=(\Delta_{1,n},\dots,\Delta_{m,n})^{\prime}$, so $\hat{V}_{n}^{\rm F}$ is also an upward-biased estimator for $\operatorname{Var}[\hat{\Delta}_{n}]$. Moreover, under mild assumptions, it is straightforward to establish using Chebyshev’s inequality that $\hat{V}_{n}^{\rm IM}$ and $\hat{V}_{n}^{\rm F}$ are consistent for their expectations in the sense of (4), and thus can be used to construct valid confidence intervals in the sense of (5).
@@
In fact, it is possible to achieve even tighter bounds by further exploiting the structure of $S_{j,\Delta}^{2}$; see Aronow et al. (2014) for details. We note, however, that we are not aware of estimators based on lower bounds other than $S_{j,\Delta}^{2}\geq 0$ which are necessarily guaranteed to be upward-biased, and in general additional assumptions need to be imposed in order to guarantee that the resulting estimators are at least asymptotically upward-biased; see, for instance, Corollary 1 in Aronow et al. (2014). As a consequence, these improved estimators may not lead to valid confidence intervals whenever these additional assumptions fail to hold.
@@ 3.3 A Novel Paired-Strata Estimator of $\operatorname{Var}[\hat{\Delta}_{n}]$
Theorem 3.2 justifies the use of $\hat{V}_{n}$ for inference about $\Delta_{n}$ and further demonstrates that the magnitude of its bias depends on the differences of the stratum-level average treatment effects across pairs of strata, and hence the bias decreases whenever the pairs of strata are formed in a way that increases their homogeneity in this sense. The bias will therefore be small not only in the extreme instance in which the stratum-level average treatment effects are similar across all strata, but also when stratum-level average treatment effects vary across strata and are only similar for adjacent strata. The latter scenario is what one might naturally expect when strata are formed by grouping individuals with similar values of $X_{i}$. In contrast, the biases of $\hat{V}^{\rm IM}_{n}$ and $\hat{V}_{n}^{\rm F}$ do not share this property. We explore this more formally in Section 4 through further restrictions on $W^{(n)}$ and $\Lambda_{n}$ that are motivated by a particular limiting thought experiment.
@@
An alternative construction for an estimator of $\operatorname{Var}[\hat{\Delta}_{n}]$ based on our paired strata $\{(\lambda_{2j-1},\lambda_{2j}):1\leq j\leq\lfloor m/2\rfloor\}$ would be to apply the estimation strategy described in (9) to the paired strata directly. However, such an estimator is not guaranteed to be upward-biased in general. In particular, it can be shown that in the case of a matched pairs design (i.e. when $\ell=1,k=2$) that this estimator coincides with the estimator $\hat{V}^{\rm alt}_{n}$ proposed in Section 5. There we demonstrate via simulation that confidence intervals for $\Delta_{n}$ constructed using this estimator may not have correct coverage in general.
Remark 3.3.
Because the bias of $\hat{V}_{n}^{\rm IM}$ only depends $\Delta_{j,n}-\Delta_{n}$, it may be minimized in extreme instances when strata are formed quite unconventionally. To illustrate this counterintuitive phenomenon, suppose for example $\eta=\frac{1}{2}$, $n$ is even, and $X_{i}=\frac{i}{n}$ for $1\leq i\leq n$. Further suppose both $Y(1)$ and $Y(0)$ are linear functions of $X$. Suppose the strata are formed as $\lambda_{j}=\{j,n+1-j\}$ for $1\leq j\leq\frac{n}{2}$, i.e., the unit with the smallest value of $X$ is matched with the unit with the largest value of $X$, the second smallest with the second largest, and so on. Then, the linearity of potential outcomes implies $\Delta_{j,n}\equiv\Delta_{n}$, so that the bias in (11) is 0. On the other hand, a natural way to form the strata, and one that minimizes the sum of squared pairwise distances of $X$, is given by $\lambda_{j}=\{2j-1,2j\}$ for $1\leq j\leq\frac{n}{2}$, under which the bias in (11) is not 0.
@@ 4 Comparison with Other Variance Estimators
In this section, we provide a framework that permits us to further discriminate among the different variance estimators of $\operatorname{Var}[\hat{\Delta}_{n}]$ discussed in Section 3. This framework is formalized by restrictions on the sequence of finite populations and experimental designs $(W^{(n)},\Lambda_{n})$ specified in the assumptions below. These restrictions are motivated by a limiting thought experiment in which $W^{(n)}$ is itself a realization of an i.i.d. sample from a fixed probability distribution, and the strata $\Lambda_{n}$ are formed in such a way that units with similar observable characteristics are grouped together, but, as explained in Remark 4.1 below, the restrictions can also be shown to hold in other contexts as well.
@@
For the analysis of $\hat{V}_{n}^{\rm F}$ only, we need to further impose the following assumption:
@@
    For any $d\in\{0,1\}$ and sequence $\{(i_{j},i_{j}^{\prime}):(i_{j},i_{j}^{\prime})\in\lambda_{j},i_{j}\neq i_{j}^{\prime},1\leq j\leq m\}$, as $n\rightarrow\infty$,
@@
Another setting in which Assumptions 4.14.2 are naturally satisfied is as follows. Let $X_{i}=\frac{i}{n}$ for $1\leq i\leq n$, so that $X^{(n)}$ are equi-distant in the unit interval. Further assume $Y_{i}(d)=f_{d}(X_{i})$ for $d\in\{0,1\}$, where $f_{0}$ and $f_{1}$ are deterministic Lipschitz functions. Define $Q$ such that $\tilde{X}\sim\mathrm{Unif}[0,1]$ under $Q$ and $\tilde{Y}(d)=f_{d}(\tilde{X})$ degenerately for $d\in\{0,1\}$. Then, it can be shown using elementary properties of the Riemann integral that Assumptions 4.14.2 hold.
@@
In the remainder of the section we compare the limits of $\hat{V}_{n}$, $\hat{V}^{\rm IM}_{n}$ and $\hat{V}_{n}^{\rm F}$ relative to the asymptotic variance $V$ defined above. Recall from Section 3 that it is possible to establish under weak assumptions on the sequence of population “moments” that each of these variances is consistent for their expectations in the sense of (4). Accordingly, in what follows it suffices to study the limit of the expectations of these estimators under Assumption 4.1, and compare them to $V$.

### v3

@@ Abstract
This paper considers the problem of design-based inference for the average treatment effect in finely stratified experiments. Here, by “design-based” we mean that the only source of uncertainty stems from the randomness in treatment assignment itself; by “finely stratified” we mean that units are first stratified into groups of a fixed size according to baseline covariates and then, within each group, a fixed number of units are assigned uniformly at random to treatment and the remainder to control. In this setting we present a novel estimator of the variance of the difference-in-means based on pairing “adjacent” strata. Importantly, our estimator is well defined even in the challenging setting where there is exactly one treated or control unit per stratum. We prove that our estimator is upward-biased, and thus can be used for inference under mild restrictions on the finite population. We compare our estimator with some well-known estimators that have been proposed previously in this setting, and demonstrate that, while these estimators are also upward-biased, our estimator has smaller bias and therefore leads to more precise inferences whenever “adjacent” strata are sufficiently similar. To further understand when our estimator leads to more precise inferences, we then introduce a framework motivated by a thought experiment in which the finite population can be modeled as having been drawn once in an i.i.d. fashion from a well-behaved probability distribution. In this framework, we argue that our estimator dominates the others in terms of limiting bias and that these improvements are strict except under exceptionally strong restrictions on the treatment effects. Finally, we illustrate the practical relevance of our theoretical results through a simulation study, which reveals that our estimator can in fact lead to substantially more precise inferences, especially when the quality of stratification is high.
@@ 1 Introduction
This paper considers the problem of design-based inference on the average treatment effect in finely stratified experiments. Here, by “design-based” we mean that there is no sampling uncertainty and the only randomness stems from the variation in treatment assignment itself; by “finely stratified” we mean that treatment is assigned in the following fashion: units are first stratified into groups of a fixed size $k$ according to baseline covariates and then, within each group, a fixed number $\ell<k$ are assigned uniformly at random to treatment and the remainder to control. A prominent special case of this framework is a matched pairs design, in which $k=2$ and $\ell=1$. While our primary focus is on experiments where treatment is assigned at the level of the individual, in Remark 2.1 we explain how our analysis can accommodate experiments where the treatment is assigned at the cluster level.
@@
We then compare our variance estimator with a canonical variance estimator that has been proposed previously for the case in which $\ell=1$ or $k-\ell=1$. The estimator, proposed by Imai (2008), is based on the sample variance obtained by viewing the difference-in-means estimates for each stratum as independent observations. Existing results demonstrate that this estimator is also upward-biased for the variance of the difference-in-means estimator, but we demonstrate that the bias of our estimator is lower (and as a consequence leads to more precise inferences) whenever we can ensure that the “adjacent” strata used to construct our estimator are sufficiently similar. To further distinguish our estimator from the estimator proposed in Imai (2008), as well as a second estimator proposed by Fogarty (2018a) which modifies the Imai estimator by partialling out the average covariate values in each stratum, we introduce a framework motivated by a thought experiment in which the finite population can be modeled as being drawn once in an i.i.d. fashion from a well-behaved probability distribution. In this framework, we show that the limiting bias of our estimator is strictly smaller than that of Imai (2008) unless the treatment effects are homogeneous, and strictly smaller than that of Fogarty (2018a) unless the conditional average treatment effect is linear in the covariates. As a result, confidence intervals based on our variance estimator will be strictly shorter than those based on the other variance estimators unless these exceptional restrictions hold.
@@ 2 Setup and Notation
$\hat{\Delta}_{n}^{\dagger}=\frac{1}{n(1)}\sum_{1\leq i\leq n}Y_{i}^{\dagger}D_{i}-\frac{1}{n(0)}\sum_{1\leq i\leq n}Y_{i}^{\dagger}(1-D_{i})~{},$
@@ 3.1 Motivating Result
When (7) holds, we say that $\tilde{V}_{n}$ is upward-biased for $\operatorname{Var}[\hat{\Delta}_{n}]$. Theorem 3.1 thus demonstrates that, under mild regularity conditions, confidence intervals for $\Delta_{n}$ of the form $[\hat{\Delta}_{n}\pm\sqrt{\tilde{V}_{n}}\cdot z_{1-\alpha/2}]$ are valid in the sense of having the right limiting coverage probability, if and only if $\tilde{V}_{n}$ is asymptotically upward-biased for $\operatorname{Var}[\hat{\Delta}_{n}]$. These assumptions include weak restrictions on the finite population (i.e., (2)), a non-degeneracy condition (i.e., (3)), and a requirement that $n\cdot\tilde{V}_{n}$ is consistent for its expectation (i.e., (4)). This last condition will typically hold under some further restrictions on finite population moments; see, e.g., Theorem 4.2 below. Finally, we note that it is immediate from this result that the length of the interval $[\hat{\Delta}_{n}\pm\sqrt{\tilde{V}_{n}}\cdot z_{1-\alpha/2}]$ will necessarily be (asymptotically, weakly) shorter whenever it is constructed using a variance estimator $\tilde{V}_{n}$ with correspondingly smaller asymptotic upwards bias. As a consequence, variance estimators with smaller bias lead to (asymptotically) more precise inferences.
@@ 3.2 Review of Some Upwards Biased Estimators of $\operatorname{Var}[\hat{\Delta}_{n}]$
is the difference in means in the $j$th stratum. This variance estimator serves as the basic scaffolding for many of the recent estimators proposed in the literature on inference in stratified randomized experiments: it is a special case of the small-block variance estimator proposed in Pashley and Miratrix (2021) when all strata are the same size (see also Zhu et al. (2024)), a special case of the pair-cluster variance estimator of De Chaisemartin and Ramirez-Cuellar (2024) in the setting of an individual-level randomized experiment, and a special case of the regression-based estimator due to Fogarty (2018a), which we introduce in Section 4. Note that the bias of $\hat{V}^{\rm IM}_{n}$ is given by (see Imbens and Rubin, 2015; Fogarty, 2018a)
@@
so that $\hat{V}^{\rm IM}_{n}$ is an upward-biased estimator of $\operatorname{Var}[\hat{\Delta}_{n}]$. Moreover, under mild assumptions, it is straightforward to establish using Chebyshev’s inequality that $\hat{V}_{n}^{\rm IM}$ is consistent for its expectation in the sense of (4), and thus can be used to construct valid confidence intervals in the sense of (5).
@@
In fact, it is sometimes possible to achieve even tighter bounds; see Aronow et al. (2014) for details. We note, however, that we are not aware of estimators based on lower bounds other than $S_{j,\Delta}^{2}\geq 0$ which are necessarily guaranteed to be upward-biased, and in general additional assumptions need to be imposed in order to guarantee that the resulting estimators are at least asymptotically upward-biased; see, for instance, Corollary 1 in Aronow et al. (2014). As a consequence, these improved estimators may not lead to valid confidence intervals whenever these additional assumptions fail to hold.
@@ 3.3 A Novel Paired-Strata Estimator of $\operatorname{Var}[\hat{\Delta}_{n}]$
Theorem 3.2 justifies the use of $\hat{V}_{n}$ for inference about $\Delta_{n}$ and further demonstrates that the magnitude of its bias depends on the differences of the stratum-level average treatment effects across pairs of strata, and hence the bias decreases whenever the pairs of strata are formed in a way that increases their homogeneity in this sense. The bias will therefore be small not only in the extreme instance in which the stratum-level average treatment effects are similar across all strata, but also when stratum-level average treatment effects vary across strata and are only similar for adjacent strata. Corollary 3.1 documents the relative biases of $\hat{V}_{n}$ versus $\hat{V}^{\rm IM}_{n}$.
Corollary 3.1.
Suppose $D^{(n)}$ satisfies Assumption 2.1. Then
$E[\hat{V}_{n}]-\operatorname{Var}[\hat{\Delta}_{n}]\leq E[\hat{V}^{\rm IM}_{n}]-\operatorname{Var}[\hat{\Delta}_{n}]$
if an only if
$\frac{1}{m}\sum_{1\leq j\leq\lfloor\frac{m}{2}\rfloor}(\Delta_{2j-1,n}-\Delta_{n})(\Delta_{2j,n}-\Delta_{n})\geq-\frac{1}{2m(m-1)}\sum_{1\leq j\leq m}(\Delta_{j,n}-\Delta_{n})^{2}~{}.$ (12)
We note that, a sufficient condition for (12) to hold is that the left-hand side is non-negative: we expect this to be the case whenever “similar” strata are paired together. We argue that such a condition holds quite generally for large $m$ in Section 4 through additional restrictions on $W^{(n)}$ and $\Lambda_{n}$ that are motivated by a particular limiting thought experiment.
@@
An alternative construction for an estimator of $\operatorname{Var}[\hat{\Delta}_{n}]$ based on our paired strata $\{(\lambda_{2j-1},\lambda_{2j}):1\leq j\leq\lfloor m/2\rfloor\}$ would be to apply the estimation strategy described in (9) to the paired strata directly. However, such an estimator is not guaranteed to be upward-biased in general. In particular, it can be shown that in the case of a matched pairs design (i.e., when $\ell=1,k=2$) that this estimator coincides with the estimator $\hat{V}^{\rm alt}_{n}$ proposed in Section 5. There we demonstrate via simulation that confidence intervals for $\Delta_{n}$ constructed using this estimator may not have correct coverage in general.
@@ 4 Comparison with Other Variance Estimators
In this section, we provide a framework that permits us to further discriminate among the variance estimators of $\operatorname{Var}[\hat{\Delta}_{n}]$ discussed in Section 3. This framework is formalized by restrictions on the sequence of finite populations and experimental designs $(W^{(n)},\Lambda_{n})$ specified in the assumptions below. These restrictions are motivated by a limiting thought experiment in which $W^{(n)}$ is itself a realization of an i.i.d. sample from a fixed probability distribution, and the strata $\Lambda_{n}$ are formed in such a way that units with similar observable characteristics are grouped together, but, as explained in Remark 4.1 below, the restrictions can also be shown to hold in other contexts as well.
@@
For the analysis of the estimator proposed by Fogarty (2018a), introduced below, we need to further impose the following assumption:
@@
    For any $d\in\{0,1\}$ and sequence $\{(i_{j},i_{j}^{\prime}):(i_{j},i_{j}^{\prime})\in\lambda_{j}^{2},i_{j}\neq i_{j}^{\prime},1\leq j\leq m\}$, as $n\rightarrow\infty$,
@@
Another setting in which Assumptions 4.14.2 are naturally satisfied is as follows. Let $X_{i}=\frac{i}{n}$ for $1\leq i\leq n$, so that $X^{(n)}$ are equi-distant in the unit interval. Further assume $Y_{i}(d)=f_{d}(X_{i})$ for $d\in\{0,1\}$, where $f_{0}$ and $f_{1}$ are deterministic Lipschitz functions. Define $Q$ such that $\tilde{X}\sim\mathrm{Unif}[0,1]$ under $Q$ and $\tilde{Y}(d)=f_{d}(\tilde{X})$ degenerately for $d\in\{0,1\}$. Then, it can be shown using elementary properties of the Riemann integral that Assumptions 4.14.2 hold. More generally, an alternative limiting thought experiment that would yield similar results allows $W^{(n)}$ to be a realization of a sample taken without replacement from a finite population that is converging (in distribution) to $Q$.
@@
In the remainder of the section we compare the asymptotic variance $V$ to the (appropriately scaled) limit of $\hat{V}_{n}$ and to the limits of the variance estimators proposed in Imai (2008) (i.e., $\hat{V}^{\rm IM}_{n}$), as well as the variance estimator proposed by Fogarty (2018a), which we now describe. Let $R$ be an $m\times L$ matrix for $L<m$ with rank $L$ and $H_{R}=R(R^{\prime}R)^{-1}R^{\prime}$ denote the projection matrix onto its column space. For each $R$, consider the estimator given by
for $\bar{X}_{j,n}=\frac{1}{k}\sum_{i\in\lambda_{j}}X_{i}$ and $\mu_{X,n}=\frac{1}{n}\sum_{1\leq i\leq n}X_{i}$. In particular, Fogarty (2018a) argues that $\hat{V}_{n}^{\rm F}$ can be less conservative than $\hat{V}_{n}^{\rm IM}$ whenever the covariates are predictive of the treatment effects in an appropriate sense. The bias for $\hat{V}_{n}^{\rm F}$ is (see Fogarty, 2018a)
where $\delta_{n}=(\Delta_{1,n},\dots,\Delta_{m,n})^{\prime}$, so $\hat{V}_{n}^{\rm F}$ is also an upward-biased estimator for $\operatorname{Var}[\hat{\Delta}_{n}]$.
As discussed in Section 3, it is possible to establish under weak assumptions on the sequence of population “moments” that each of these variances is consistent for their expectations in the sense of (4). Accordingly, in what follows it suffices to study the limit of the expectations of these estimators under Assumption 4.1, and compare them to $V$.